# No Size Fits All: The Perils and Pitfalls of Leveraging LLMs Vary with Company Size

### v1

@@ Abstract
Large language models (LLMs) are playing a pivotal role in deploying strategic use cases across a range of organizations, from large pan-continental companies to emerging startups. The issues and challenges involved in the successful utilization of LLMs can vary significantly depending on the size of the organization. It is important to study and discuss these pertinent issues of LLM adaptation with a focus on the scale of the industrial concerns and brainstorm possible solutions and prospective directions. Such a study has not been prominently featured in the current research literature. In this study, we adopt a threefold strategy: first, we conduct a case study with industry practitioners to formulate the key research questions; second, we examine existing industrial publications to address these questions; and finally, we provide a practical guide for industries to utilize LLMs more efficiently.
@@ 1 Introduction
To this end, we conduct a study with a threefold strategy, firstly, we conduct a rigorous case study of real-world practitioners from the IT industry, who are trying to work on AI adaptation and formulate three guiding research questions. RQ1. How have industries adopted LLMs so far, and what challenges do they face? RQ2. What are the barriers hindering the full utilization of LLMs in industrial applications, and how can these barriers be overcome? RQ3. How can various industries advance to maximize the utility of LLMs in practical applications? Subsequently, with an aim to address guiding research questions, we perform a thorough scoping survey of existing research publications from industrial entities of all sizes. Finally, we discuss our takeaways and insights and present a practical pilot scenario-based guide for industries to adapt to LLMs in a more informed manner.
@@ 3 Methodology
This section aims to understand, how have industries adopted LLMs so far, and what challenges they face? (RQ1).
@@ 3.1 Industrial Case Study on LLMs
We conduct an industrial case study to understand, how the LLMs are shaping industry practices, identifying both challenges and benefits. Through a meticulous process of expert consultation and iterative refinement, the questionnaire was designed to capture insightful data and serve as a tool for understanding the evolving role of LLMs in the industry. This case study covers a multitude of aspects related to LLM usage for specific application domains, corresponding risks, trust attributes, and challenges. In crafting a succinct questionnaire, our objective was to gauge the adoption and impact of LLMs in various industries. These questions can be found in Appendix B Table 4. We receive 26 responses in total from real-world practitioners of the IT industry.
@@ 3.2 Quantitative Analysis
Participants of the case study. We shared the questionnaire with the IT professionals, who are either working on LLMs or developed some solutions. The participants are industry professionals and practitioners ranging in their expertise from beginner to expert level.
Widely adapted applications by leveraging LLMs. Even though LLMs are being utilized for various applications, we observe that the majority of the industrial applications are related to financial, retail, security, and healthcare domains.
@@
Widely used LLMs. Our case study suggests that more than 50% of the applications utilize the GPT-3.5 and GPT-4 models. Recently, researchers have been assessing the capabilities of LLaMA-2 Touvron et al. (2023) and Mistral Jiang et al. (2023a).
Prompting strategy. We observe that zero-shot, few-shot, and in-context learning prompting strategies are widely adapted compared to Fine-tuning.
@@
Trust attributes to be considered. We observe that robustness, security, and hallucination are the major attributes that need to be considered to utilize the LLMs as shown in Figure 1.c.
Moreover, to gain a better understanding of the barriers to leverage the LLMs for industrial use cases, we survey 68 research papers specifically from the industry. With this study, we compile several prominent challenges and collate potential solutions to address them. The selection criteria for the papers can be referred to in the Appendix A.
@@ 4.1.1 Pre-training data issues
Potential privacy risks. To deploy large language models (LLMs) on cloud platforms, robust data privacy protocols are required to handle extensive sensitive datasets while pre-training. Key challenges include mitigating data breaches and preventing unauthorized extraction of sensitive information. Despite the adoption of LLMs in applications like disaster response management Goecks and Waytowich (2023), public health intervention Jo et al. (2023), and assisting AAC users Valencia et al. (2023), there is a noticeable lack of focus on privacy and security aspects. Moreover, it is imperative to address the potential risks associated with deploying LLMs in high-stakes scenarios.
Regulations. GDPR in Europe and CCPA in California introduce stringent guidelines for deploying LLMs by enforcing strict data handling and intellectual property rules for transparency and fairness. As highlighted by Mesko and Topol (2023), adhering to these laws in sensitive domains like healthcare is crucial to avoid harm and protect privacy.
@@ 4.1.2 Usage of APIs
Potential solutions. 1. Implementing robust security and privacy techniques, such as federated learning, is essential to safeguard user data while maintaining the functionality of LLMs, 2. A strategic way of crafting prompts is essential to avoid PII leakage Kim et al. (2023).
@@ 4.2 Reliability of LLMs’ Responses
Control the level of AI proactivity. LLMs should reduce social awkwardness, improve expressiveness, and adapt to different scenarios Liu et al. (2023c); Urlana et al. (2023). The open-ended generation of LLMs makes it challenging to customize dialog systems for public health intervention applications Jo et al. (2023).
@@
Potential solutions. Techniques such as Retrieval-Augmented Generation (RAG) are effective in reducing hallucinations. However, RAG systems struggle with complex questions requiring additional information and generate out-of-context content. Moreover, techniques such as diverse beam search Vijayakumar et al. (2016), confident decoding Tian et al. (2019) are promising in mitigating hallucinations. Additionally, model editing techniques Hoelscher-Obermaier et al. (2023) can address the unintended associations, enhancing the practical usage of LLMs.
@@ 4.3 Infrastructure Accessibility
Carbon emissions. Infrastructure is crucial for deploying LLMs, influencing factors like processing speed, latency, cost, and training needs. High-performance hardware is necessary to boost speed and reduce latency, enhancing user experience but demanding careful budgeting due to high costs. Achieving an optimal balance between cost and performance is crucial for the efficient training and scalability of LLM applications.
@@
Potential Solution. Open-source models like LLaMA-2 Touvron et al. (2023) are more favorable for industrial deployment. Further studies on efficient model optimization techniques such as quantization, pruning, and distillation are required Laskar et al. (2023). Moreover, closed-source models that can utilize prompt compression techniques such as LLMLingua Jiang et al. (2023b) are promising.
@@ 4.4 Domain Adaption
Lack of domain-specific datasets. The ability of LLMs in the finance and medical domains is lacking due to insufficient Liu et al. (2023a) domain-specific training data in the foundation models Li et al. (2023c). Consequently, the current versions of GPT-4 and ChatGPT do not meet the industrial requirements to build financial analyst agents Li et al. (2023c). While LLMs can generate relevant reasoning, they fall short of the desired standard, indicating significant room for improvement.
Lack of expert domain data. Majority of the LLMs struggle to perform the expert data domain tasks Alfassy et al. (2022), such as image-text and text-image retrieval in technical documentation.
Diversity. Currently, LLMs fail to mitigate social bias due to the absence of diverse demographic groups Lee et al. (2023) data in the training data. Foundation models should give equal importance to aspects such as ethnicity, nationality, gender, and religion. At present, most LLMs reflect the collective intelligence of Western society.
@@
Potential Solution. 1. Pre-training data should consist of various domain mixtures; however, finding the right mixture is an open challenge. 2. LLMs should be carefully tested to ensure they treat marginalized individuals and communities equally Kotek et al. (2023). 3. Continuous pre-training can help overcome the drawbacks of the in-context learning strategy.
@@ 4.5 Data Creation Using LLMs
Few works attempt to generate synthetic datasets by utilizing LLMs. However, four major concerns exist with using LLMs for synthetic data creation/annotation; 1). Lack of diversity. Synthetic datasets may lack diversity due to the limited knowledge base Ramakrishna et al. (2023) of LLMs, 2). Quality and resource requirements. The quality of the annotated data might increase with the size of the LLM used for the annotation Sun et al. (2023). However, leveraging large LLMs requires higher computational resources, 3). In-context learning (ICL) challenges. ICL is a widely adopted approach for textual task data annotation tasks Li et al. (2023c). However, the challenge lies in responsibly incorporating the model’s output to deliver value to users without misleading them or inadvertently amplifying malicious behavior Deng et al. (2023), 4). Performance of complex tasks. The performance of models degrades when using synthetic data for complex tasks.
@@ 4.6 Sub-standard Performance of LLMs
Multilingual. Maximizing LLMs’ potential in healthcare, retail, and finance requires high-quality, labeled datasets. Additionally, with most LLMs being limited to English, there is significant room for creating robust multilingual models.
Multi-modal. With the rising demand for multi-modal applications Feng et al. (2023); Lu et al. (2023b), only a few studies focused on utilizing LLMs for such industrial applications. More efforts are needed to integrate LLMs with voice assistants and Robotics Yamazaki et al. (2023).
@@ 4.7 Evaluation of LLMs
In sectors like legal, finance, and healthcare, blending LLMs with human feedback is crucial to lowering false positives, underscoring the importance of human oversight in safety-critical applications Liu et al. (2023b). Moreover, our analysis (see Appendix C) reveals that less than 15% of studies conduct human evaluations to assess LLM outputs, indicating a need for more rigorous validation methods. Evaluating long-form question answering is challenging for LLMs Zhao et al. (2023), as additional contextual information may not always be available in practical QA scenarios. Current metrics like BLEU Papineni et al. (2002) and ROUGE Lin (2004) primarily evaluate the explanation task, but these measures are insufficient for assessing the reasonableness of LLM responses.
@@ 4.8 Ethical Concerns
Building trust. LLMs face gender bias issues while generating professional documents like recommendation letters Li et al. (2023c). Further, most of the studies are limited to understanding gender bias in male and female pronouns, indicating a need for more research on marginalized groups, such as the LGBTQIA+ community. Moreover, AI-assisted writing should be employed judiciously to prevent reinforcing gender stereotypes Wan et al. (2023). Increasing trustworthiness involves prioritizing the addressing, opposing, and limiting of further proliferation of bias Kotek et al. (2023). Recent studies also suggest to conduct experiments with LLMs’ at least three times to validate the results Chen et al. (2023).
Enhancing creativity. The involvement of AI models should not aim to replace the author’s thinking but rather enhance it by producing new ideas and discover connections or findings Shen et al. (2023).
Fairness in data visualization. Developing interactive data visualization techniques/frameworks to identify the hidden biases and ensure fairness Kwon and Mihindukulasooriya (2023).
Linking models. Techniques such as LLM Attribution Foley et al. (2023) are promising for linking fine-tuned models with their pre-trained versions.
Protecting integrity. To maintain the integrity of LLMs, implementing guardrails such as NeMo Rebedea et al. (2023), LangKit [^1] https://docs.whylabs.ai/docs/langkit-api/, and TrustLLM Sun et al. (2024) helps to prevent the generation of biased, offensive, or inaccurate responses.
Addressing these challenges requires a combination of technical expertise, ethical considerations, and ongoing research efforts. As shown in Figure 2, we categorize each paper (total of 68) based on its application life cycle and observed that, due to the above-mentioned pitfalls, more than 70% of LLM-based studies are still in the conceptual phase.
@@ 5 Maximizing LLM Utilization Across Industries
This section offers a suggestive guide to various industries to maximize the utilization of LLMs for Natural Language Generation (NLG) applications (RQ3). As shown in Table 1, our suggestions are tailored to various industries, considering their distinct goals, resources, and workforce capabilities. Furthermore, the recommendations intended for small and medium-sized industries equally apply to large-scale industries.
1) Small-scale industries such as startups with less than 100 employees need to optimize the use of LLMs within constraints of limited computational resources and workforce. These industries should emphasize prompt engineering and transfer learning techniques to utilize robust small LLMs with up to 3 billion parameters with permissive licenses. Further, these industries should focus on monolingual uni-modal tasks and actively perform the inference on a few hundred samples. To reduce the inference duration, these industries should opt for optimization techniques such as quantization. Moreover, these industries encounter challenges such as potential reductions in model accuracy, cost considerations, and the need for technical expertise, some of these can be addressed by partnering with AI consulting firms.
@@
3) Large scale industries such as MNCs should focus on continuous pre-training of LLMs while ensuring compliance with regulatory requirements. These industries can leverage LLMs effectively across multi-lingual, cross-lingual, and multi-modal generation tasks. Training of such models can take from a few weeks to months, which requires high-quality data and huge compute as well. These industries should focus on establishing several collaborative tools and frameworks to maximize LLM utilization.
For all industries, we recommend using open-source models with proper licenses to address ethical concerns and comply with LLM regulatory guidelines. Additionally, robust testing and validation protocols are essential to meet industry standards. Fostering strong collaborations and knowledge sharing between industry and academia is crucial for advancing responsible LLM development and deployment.
@@ 6 Conclusions
This study delves into the utilization of Large Language Models (LLMs) through an industrial lens, with a specific focus on identifying roadblocks to their adoption. It meticulously examines various pitfalls and provides potential solutions. Moreover, this study offers a guide to organizations of all sizes to maximize the utilization of LLMs for industrial use cases. By identifying pitfalls and suggesting potential directions, the study offers a strategic roadmap for optimizing LLM effectiveness in industrial operations. We plan to maintain a GitHub repository with the most recent papers in the field.

### v2

@@ Abstract
Large language models (LLMs) are playing a pivotal role in deploying strategic use cases across a range of organizations, from large pan-continental companies to emerging startups. The issues and challenges involved in the successful utilization of LLMs can vary significantly depending on the size of the organization. It is important to study and discuss these pertinent issues of LLM adaptation with a focus on the scale of the industrial concerns and brainstorm possible solutions and prospective directions. Such a study has not been prominently featured in the current research literature. In this study, we adopt a threefold strategy: first, we conduct a case study with industry practitioners to formulate the key research questions; second, we examine existing industrial publications to address these questions; and finally, we provide a practical guide for industries to utilize LLMs more efficiently. We release the GitHub [^1] https://github.com/vinayakcse/IndustrialLLMsPapers repository with the most recent papers in the field.
@@ 1 Introduction
To this end, we conduct a study with a threefold strategy, firstly, we conduct a rigorous case study of real-world practitioners from the IT industry, who are trying to work on AI adaptation and formulate three guiding research questions. RQ1. How have industries adopted LLMs so far, and what challenges do they face? RQ2. What are the barriers hindering the full utilization of LLMs in industrial applications, and how can these barriers be addressed? RQ3. How can various industries advance to maximize the utility of LLMs in practical applications? Subsequently, with an aim to address guiding research questions, we perform a thorough scoping survey of existing research publications from industrial entities of all sizes. Finally, we discuss our takeaways and insights and present a practical pilot scenario-based guide for industries to adapt to LLMs in a more informed manner.
@@ 3 Methodology
This section aims to explore how industries have adopted LLMs and the challenges they face (RQ1).
@@ 3.1 Industrial Case Study on LLMs
We conduct an industrial case study to understand, how the LLMs are shaping industry practices, identify the underlying challenges and benefits. Through a meticulous process of expert consultation and iterative refinement, the questionnaire was designed to capture insightful data and serve as a tool for understanding the evolving role of LLMs in the industry. This case study covers a multitude of aspects related to LLM usage for specific application domains, corresponding risks, trust attributes, and challenges. In crafting a succinct questionnaire, our objective was to gauge the adoption and impact of LLMs in various industries. These questions can be found in Appendix B Table 4. We receive 26 responses in total from real-world practitioners of the IT industry. We did a case study on 26 companies which are leveraging LLMs for their use-cases. This exercise is non-trivial as most companies have not made their LLM-related use cases public.
@@ 3.2 Quantitative Analysis
Participants of the case study. We shared the questionnaire with the IT professionals, who are either working on LLMs or have developed some solutions. The participants are industry professionals and practitioners with expertise ranging from beginner to expert level.
Widely adapted applications by leveraging LLMs. Even though LLMs are being utilized for various applications, we observe that the majority of these industrial applications are related to financial, retail, security, and healthcare domains.
@@
Widely used LLMs. Our case study indicates more than 50% of the applications utilize the GPT-3.5 and GPT-4 models. Recently, researchers have been assessing the capabilities of LLaMA-2 Touvron et al. (2023) and Mistral Jiang et al. (2023a).
Prompting strategy. We observe that zero-shot and in-context learning prompting strategies are widely adapted compared to fine-tuning.
@@
Trust attributes to be considered. We observe that robustness, security, and hallucination are the major challenges that need to be considered to utilize the LLMs as shown in Figure 1.c.
Moreover, to gain a better understanding of the barriers to leverage the LLMs for industrial use cases, we also survey 68 research papers specifically from the industry. In this study, we compile several prominent challenges and present potential solutions to address them. The selection criteria for the papers can be found in the Appendix A.
@@ 4.1.1 Pre-training data issues
Potential privacy risks. To deploy large language models (LLMs) on cloud platforms, robust data privacy protocols are required to handle extensive sensitive datasets while pre-training. Key challenges include mitigating data breaches and preventing unauthorized extraction of sensitive information. Despite the adoption of LLMs in applications like disaster response management Goecks and Waytowich (2023), public health intervention Jo et al. (2023), and assisting Augmentative and Alternative Communication (AAC) users Valencia et al. (2023), there is noticeable lack of focus on privacy and security aspects. Moreover, it is imperative that potential risks associated with deploying LLMs in high-stakes scenarios are addressed.
Regulations. GDPR in Europe and CCPA in California introduce stringent guidelines for deploying LLMs by enforcing strict data handling and intellectual property rules to ensure transparency and fairness. As highlighted by Mesko and Topol (2023), adhering to these laws in sensitive domains like healthcare is crucial to avoid harm and protect privacy.
@@ 4.1.2 Usage of APIs
Potential solutions. 1. Robust security and privacy techniques like federated learning are essential to safeguard user data while maintaining the functionality of LLMs, 2. A strategic way of crafting prompts is essential to avoid Personally Identifiable Information (PII) leakage Kim et al. (2024).
@@ 4.2 Reliability of LLMs’ Responses
Control the level of AI proactivity. LLMs should minimize social awkwardness, enhance expressiveness, and adapt to different scenarios Liu et al. (2023b); Urlana et al. (2024). The open-ended generation of LLMs makes it challenging to customize dialog systems for public health intervention applications Jo et al. (2023).
@@
Potential solutions. Techniques such as Retrieval-Augmented Generation are effective in reducing hallucinations. However, such systems struggle with complex questions that require additional information often generating out-of-context content. Moreover, techniques such as diverse beam search Vijayakumar et al. (2018), confident decoding Tian et al. (2019) are promising in mitigating hallucinations. Additionally, model editing techniques Hoelscher-Obermaier et al. (2023) can address the unintended associations, enhancing the practical usage of LLMs.
@@ 4.3 Infrastructure Accessibility
Carbon emissions. Infrastructure is crucial for deploying LLMs, influencing factors like processing speed, latency, cost, and training needs. High-performance hardware is necessary to boost speed and reduce latency, enhancing user experience but it requires careful budgeting due to associated high costs. Achieving an optimal balance between cost and performance is crucial for the efficient training and scalability of LLM applications.
@@
Potential Solution. Open-source models like LLaMA-2 Touvron et al. (2023) are more favorable for industrial deployment. Further studies on efficient model optimization techniques such as quantization, pruning, and distillation are required Laskar et al. (2023). Moreover, closed-source models that can utilize prompt compression techniques such as LLMLingua Jiang et al. (2023b).
@@ 4.4 Domain Adaption
Lack of domain-specific datasets. The ability of LLMs in the finance and medical domains is lacking due to insufficient domain-specific training data in the foundation models Liu et al. (2024); Li et al. (2023c). Consequently, the current versions of GPT-4 and ChatGPT do not meet the industrial requirements to build financial analyst agents Li et al. (2023c). While LLMs can generate relevant reasoning, they fall short of the desired standard, indicating significant room for improvement.
Diversity. LLMs fail to mitigate social bias due to a lack of diverse demographic data Lee et al. (2023). Foundation models must equally consider factors like ethnicity, nationality, gender, and religion, as most currently reflect western perspectives.
@@
Potential Solution. 1. Pre-training data should consist of various domain mixtures; however, finding the right mixture is still an open challenge. 2. LLMs should be carefully tested to ensure they treat marginalized individuals and communities equally Kotek et al. (2023). 3. Continuous pre-training can help overcome the drawbacks of the in-context learning strategy.
@@ 4.5 Data Creation Using LLMs
Few works attempt to generate synthetic datasets by utilizing LLMs. However, three major concerns exist with using LLMs for synthetic data creation/annotation; 1). Lack of diversity. Synthetic datasets may lack diversity due to the limited knowledge base Ramakrishna et al. (2023) of LLMs, 2). Quality and compute. The quality of the annotated data might improve with the size of the LLM used for the annotation Sun et al. (2023). However, leveraging large LLMs requires higher computational resources, 3). In-context learning (ICL) challenges. ICL is a widely adopted approach for textual task data annotation tasks Li et al. (2023c). However, the main challenge lies in responsibly incorporating the model’s output is to deliver value to users without misleading them or inadvertently amplifying malicious behavior Deng et al. (2023).
@@ 4.6 Sub-standard Performance of LLMs
Multilingual and Multi-Modal: Most of the LLMs are being limited to English, there is significant room for creating robust multilingual models. Only a few studies have focused on utilizing LLMs for such multi-modal industrial applications Feng et al. (2024); Lu et al. (2023). More efforts are needed to integrate LLMs with voice assistants and Robotics Yamazaki et al. (2023).
4.7 Explainability and Interpretability
The robust performance of LLMs across various tasks underscores the importance of explainability and interpretability to foster trust in their predictions. However, several challenges impede the development of explainable models.
Black Box Nature: Many popular LLMs, such as ChatGPT and Gemini Team et al. (2023), are accessible only through APIs, limiting users’ understanding of their internal workings.
Scale and Complexity of Models: The large-scale training on vast data leads to complex models, making it hard to identify which parameters influence specific decisions Brown et al. (2020).
Performance Trade-Off: Balancing model performance with the ability to provide meaningful explanations is a significant challenge; many models struggle to maintain this equilibrium.
Language Ambiguity: The inherent ambiguity of language complicates the generation of clear explanations, as words and sentences can have multiple meanings depending on context Wang (2023).
Potential Solutions. Model Simplification: Developing simpler models can enhance interpretability, provides a clear understanding of the decision-making processes of LLMs Che et al. (2016).
Training Data Transparency: Sharing details about training datasets and their sources can illuminate knowledge gaps and potential biases in the models Bender and Friedman (2018).
Interactive Exploration Tools: Creating interactive platforms that allow users to manipulate inputs, visualize attention patterns, and observe changes in outputs can provide valuable insights into model behavior Olah et al. (2018).
@@ 4.8 Evaluation of LLMs
In sectors like legal, finance, and healthcare, blending LLMs with human feedback is crucial to lowering false positives, underscoring the importance of human oversight in safety-critical applications Liu et al. (2023a). Moreover, our analysis (see Appendix C) reveals that less than 15% of studies conduct human evaluations to assess LLM outputs, indicating a need for more rigorous validation methods. Evaluating long-form question answering is challenging for LLMs Zhao et al. (2023), as additional contextual information may not always be available in practical QA scenarios. Current metrics like BLEU Papineni et al. (2002) and ROUGE Lin (2004) primarily evaluate the similarity, but are insufficient for assessing the reasonableness of LLM responses.
@@ 4.9 Ethical Concerns
Enhancing creativity. AI models should enhance, not replace, human creativity by generating new ideas and insights Shen et al. (2023).
Fairness in data visualization. Interactive data visualization can help detect and address hidden biases Kwon and Mihindukulasooriya (2023).
Linking models. Techniques such as LLM Attribution Foley et al. (2023) link fine-tuned models to their pre-trained versions.
Protecting integrity. Guardrails such as NeMo Rebedea et al. (2023), LangKit [^2] https://docs.whylabs.ai/docs/langkit-api/, and TrustLLM Sun et al. (2024) help to maintain LLM integrity by preventing biased or inaccurate outputs.
Addressing these challenges requires a combination of technical expertise, ethical considerations, and further research efforts. In Figure 2, we categorize each paper (total of 68) based on its application life cycle and observed that, due to the above-mentioned pitfalls, more than 70% of LLM-based studies are still in the conceptual phase.
@@ 5 Maximizing LLM Utilization Across Industries
This section offers a suggestive guide to various industries to maximize the utilization of LLMs for Natural Language Generation (NLG) applications (RQ3). As shown in Table 1, our suggestions are tailored to various industries, considering their distinct goals, resources, and workforce capabilities. The recommendations for small and medium-sized industries equally apply to large-scale industries.
1) Small-scale industries such as startups with less than 100 employees need to optimize the use of LLMs within constraints of limited computational resources and workforce. These industries should emphasize prompt engineering and transfer learning techniques to utilize robust small LLMs with up to 3 billion parameters with permissive licenses. Further, these industries should focus on monolingual tasks and actively perform the inference on a few hundred samples. To reduce the inference duration, these industries should opt for optimization techniques such as quantization. Moreover, these industries encounter challenges such as potential reductions in model accuracy, costs, and need for technical expertise. Some of these can be addressed by partnering with AI consulting firms.
@@
3) Large scale industries such as MNCs should focus on continuous pre-training of LLMs while ensuring compliance with regulatory requirements. These industries can leverage LLMs effectively across multi-lingual, cross-lingual, and multi-modal generation tasks. Training such models can take from a few weeks to months, which requires high-quality data and huge compute as well. These industries should focus on establishing several collaborative tools and frameworks to maximize LLM utilization. For all industries, we recommend using open-source models with appropriate licenses to address ethical concerns and comply with LLM regulatory guidelines. Additionally, robust testing and validation protocols are essential to meet industry standards. Fostering strong collaborations and knowledge sharing between industry and academia is crucial for advancing responsible LLM development and deployment.
@@ 6 Conclusions
This study delves into the utilization of Large Language Models (LLMs) through an industrial lens, with a specific focus on identifying roadblocks to their adoption. It meticulously examines various pitfalls and provides potential solutions. Moreover, this study offers a guide to organizations of all sizes to maximize the utilization of LLMs for industrial use cases. By identifying pitfalls and suggesting potential directions, the study offers a strategic road-map for optimizing LLM effectiveness in industrial operations.